if f(x)= 2x
find f(x+3)
please help me. i dont understand functions at all,

Answer by jim_thompson5910(35256)

f(x)= 2x

f(x+3)= 2(x+3) ... replace EVERY 'x' with 'x+3'

f(x+3)= 2x+6


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim